# Benedict Quartey, Ifrah Idrees, Jason Xinyu Liu, Stefanie Tellex

### v3

@@ VI Conclusion
We introduce LAMBDA ($\lambda$), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon mobile manipulation tasks, spanning room-to-room and floor-to-floor navigation in indoor environments. Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
To evaluate the benchmark, we tested multiple models: RT-1, MotionGlot, and LIMP, to assess their performance and data efficiency on these complex tasks. Learning-based models (RT-1 and MotionGlot) demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1 and MotionGlot, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
Limitations and Future Work. While the inclusion of a quadruped like Spot in the real world tasks highlights the advantages of quadrupedal platforms, the lack of simulated quadruped data limits the ability to leverage diverse environments that simulators can provide, such as a broader range of floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enable greater generalization and provide richer datasets for evaluation. Furthermore, extending the benchmark to include tasks beyond pick-and-place, such as those involving more intricate manipulation, would increase its relevance and applicability to a wider range of robotic systems and models. Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency. Evaluating their performance on $\lambda$’s sparse rewards would offer additional insights into their suitability for long-horizon mobile manipulation tasks and broaden the scope of the benchmark.
@@
TABLE V: Ablations results. Success rates for Task Generalization on growing subsets of the simulation dataset for both RT-1 and its LSTM variant. The computation is the same as Table III. “Avg RoC” indicates the average rate of change between subset results.

### v5

@@ VI Conclusion
We introduce LAMBDA ($\lambda$), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments. Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
To evaluate the benchmark, we tested multiple models to assess their performance and data efficiency on these complex tasks. Learning-based models, RT-1 and MotionGlot, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by $\lambda$’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
Limitations and Future Work. While the inclusion of a quadruped like Spot in the real world tasks highlights the advantages of quadrupedal platforms, the lack of simulated quadruped data limits the ability to leverage diverse environments that simulators can provide, such as a broader range of floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enable greater generalization and provide richer datasets for evaluation. Furthermore, extending the benchmark to include tasks beyond pick-and-place, such as those involving more intricate manipulation, would increase its relevance and applicability to a wider range of robotic systems and models. Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency. Evaluating their performance on $\lambda$’s sparse rewards would offer additional insights into their suitability for long-horizon MoMa tasks and broaden the scope of the benchmark.
@@
TABLE V: Ablations results. Success rates for Task Generalization on growing subsets of the simulation dataset. “Avg RoC” indicates the average rate of change between subset results.